20 randomly selected students were asked if they were freshman, sophomore, junior, or senior. The following are the responses from these 20 students.

senior sophomore freshman senior junior senior freshman junior sophomore freshman junior senior junior freshman senior freshman sophomore senior junior junior

a. Construct a frequency distribution for the data.
b. Draw a histogram graph for the frequency distribution constructed in a.

c. What is the percentage of students who are not junior?

Show all calculation steps.
Answer #1

a)
Counting the categories we get the frequency distribution:

Freshman 5

Junior 6

Sophomore 3

Senior 6

Total 20

b)Histogram:


6 5 3 Freshnan Junior Sophonore Senior

c) As there are 6 juniors ut of 20 selected students. Therefore the percentage of Junior students = (6/20)*100 = 30%

Therefore the percentage of students that are not Junior = (100 - percentage of Junior students)% = (100-30)%= 70%